User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair: Ross

next week's chair: Katie

Attended: Tiege, Lucas, Patrick, Prad, Lim, Guifre, Earl



General Issues

  • Yg: Monica generating 20 GB/hour in error conditions.
  • Yg: eRemoteCtrl isn't showing right date and timing of current scan.
    • Maybe could use backup control instead, eRemote control doesn't seem to give much? Nice pie graphs? Can also use other log monitor which greps important information out.
    • Search for field system backup.
  • Ho needs to have cry pump on all night to keep temperature reasonable. Even with pump on may be around 80-90K. Not much life left in pump.

Handover Issues

  • AUA056
    • Ho/Yg: auto-correlation did not update, missing checkmk5 block in the procedure file.
      • slogit didn't run, got emailed around. Manually drudged, checkmk5 not added.
  • AOV039
    • Yg: Monica log file /usr2/oMlog oMonica.log was writing flat out at 20 GB per hour. This was due to constant errors logging that Monica couldn't connect to weather on localhost port 5000. A restart of Monica server on pcfsyg fixed the problem.
  • R4910
    • Yg: Field system had locked up; missed scans from 248-2130b.
  • R1911
    • Yg: Waiting at limit for several scans. Field system had forced antenna to limit and would not take new source commands. Restarted the field system and schedule at scan 253-1025a.
      • Due to old version of field system maybe.
/home/www/auscope/opswiki/data/pages/operations/agenda2019_09_19.txt · Last modified: 2019/09/19 03:13 by Ross Turner